# Typical infinite mixing automorphisms are rank one

### v1

@@
Typical infinite mixing automorphisms are rank one
@@
Bashtanov proved that generic mixing automorphisms of probability space with respect to the Alpern-Tikhonov metric have rank one. Using Sidon’s constructions, we showed that generic infinite mixing automorphisms also have rank one.
@@ 1 Introduction
Alpern and Tikhonov [6] introduced a complete metric on the space of mixing actions. Bashtanov [1] proved that the class of rank-one actions is generic in the space ${\bf Mix}$ of invertible mixing transformations with invariant probability measure. We will establish a similar theorem for the space ${\bf Mix_{\infty}}=\{T:\,T^{n}\to 0\}$ of transformations of $T$ with sigma-finite invariant measure. We will prove that the conjugacy class of the Sidon transformation of rank one (see [4], [5]) is dense in ${\bf Mix_{\infty}}$. Since the class of rank-one transformations is a $G_{\delta}$-set, taking into account the density of the latter, we obtain the genericity of this class. A consequence of this typicality, given the result [2], is the typicality of the minimality of the centralizer (the transformation $T$ commutes only with its powers).
@@
where $\{A_{i}\}$ is some fixed family dense in the family of all measurable sets of finite measure. The space $({\bf Aut_{\infty}},\rho)$ is a complete separable space. It is easy to show that in $({\bf Aut_{\infty}},\rho)$ the so-called mixing transformations form a set of the first category (countable intersection of nowhere dense sets). If we strengthen the Halmos metric to the Alpern-Tikhonov metric, we obtain a complete metric space ${\bf Mix}_{\infty}$). The question arises: which invariants, i.e. properties preserved under conjugation in $\bf Aut_{\infty}$) are typical? A property is called typical if a dense $G_{\delta}$-set of transformations has this property.
For example, given the known methods, it can be shown that in ${\bf Mix}_{\infty}$ the spectrum of all symmetric powers of the transformation is typically single. It easily follows from the results of [5] that a typical transformation in ${\bf Mix}_{\infty}$ is not isomorphic to its inverse.
@@
The sequence $T_{k}$, fundamental in $({\bf Mix_{\infty}},r)$, converges to some $T$ by virtue of the completeness of the Halmos metric in the space ${\bf Aut}(\mu_{\infty})$. It is easy to show that in this case $T_{k}\to T$ and in the metric $\rho$.
2 Density of the conjugacy class of a Sidon automorphism
Let us recall the necessary definitions.
Construction of rank one. Fix a natural number $h_{1}$, a sequence $r_{j}\to\infty$ (parameter $r_{j}$ is the number of columns into which the tower of stage $j$ is cut) and a sequence of integer vectors (parameters of superstructures)
$\bar{s}_{j}=(s_{j}(1),s_{j}(2),\dots,s_{j}(r_{j}-1),s_{j}(r_{j})).$
At step $j=1$, by definition, a set of half-intervals $E_{1},SE_{1},S^{2}E_{1},\dots,S^{h_{1}-1}E_{1}$ is given. Let at step $j$ a system of non-intersecting half-intervals
@@ 2 Density of the conjugacy class of a Sidon automorphism
be defined, and on the half-intervals $E_{j},SE_{j},\dots,S^{h_{j}-2}E_{j}$ the transformation $S$ is a parallel translation. Such a set of half-intervals is called the tower of step $j$, their union is denoted by $X_{j}$ and is also called a tower.
@@
Let the construction $S$ of rank one have the following property: the intersection $X_{j}\cap S^{m}X_{j}$ for $h_{j}<m\leq h_{j+1}$ can be contained only in one of the columns $X_{i,j}$ of the tower $X_{j}$. Such a construction is called Sidon. The measure of the space $X$ in this case is infinite.
Theorem 1. Let $T$ be a mixing transformation, $S$ a Sidon transformation of rank one. For every $\varepsilon>0$ there exists a conjugation $R$ such that $r(R^{-1}SR,T)<\varepsilon$.
@@
The proof of the lemma is simple, we leave it as an exercise.
The proof of the theorem is brief. Let us consider a part of the Kakutani tower:
@@
We take a very thin tower $X_{j}$, appearing in the construction of the Sidon transformation $S$, and wind on $U_{N}$ the tower $\tilde{X}_{j}$ corresponding to the conjugation $\tilde{S}=R^{-1}SR$ in such a way that the following conditions are satisfied:
@@
(2) after the moving floor of the wound tower leaves the top of $U_{N}$ for the $k$-th time, it must travel outside $U_{N}$ for a time $s_{k}>10s_{k-1}$ ($s_{1}>10N$). This method can be called the Sidon winding, which will ensure that the intersection measure is small:
@@
In reality, when choosing a large $N$ and a suitable winding, instead of equalities we get inequalities with a very small right-hand side.
So for $\varepsilon>0$ there are always $j$, $N$ and $R$ such that at the intersection $\tilde{X}_{j}\cap U_{N}$ the transformations $\tilde{S}$ and $T$ almost coincide, hence are close in the Halmos metric, but their degrees are uniformly close with respect to the $d_{w}$ metric. Thus we obtain $r(R^{-1}SR,T)<\varepsilon$ for a suitable $R$.
The set of rank-one transformations is a $G_{\delta}$ set (we use the proof from [1] without any changes). We have found that it is dense. Thus, the typicality of rank-one is proved.
In [3] we used a similar trick, but the windings were Ornstein and the spacers for flows and automorphisms of rank one were Ornstein as well. In the case of infinite measure, a pleasant opportunity arises to use large Sidon constructions, which allows us to use a simpler and more constructive method. For other results on Sidon transformations, see [4].
@@
Альперн и Тихонов [6] ввели полную метрику на пространстве перемешивающих действий. Баштанов [1] доказал, что класс действий ранга один типичен в пространстве ${\bf Mix}$братимых перемешивающих преобразований с инвариантной вероятностной мерой. Мы установим аналогичную теорему для пространства ${\bf Mix_{\infty}}=\{T:\,T^{n}\to 0\}$реобразований $T$игма-конечной инвариантной мерой. Докажем, что класс сопряженности сидоновского преобразования ранга один (см. [4], [5]) плотен в ${\bf Mix_{\infty}}$. Так как класс преобразований ранга один является $G_{\delta}$-множеством, с учетом плотности последнего получаем типичность этого класса. Следствием этой типичности с учетом результата [2] является типичность минимальности централизатора (преобразование $T$оммутирует только со своими степенями).

### v2

# Typical infinite mixing automorphisms are rank-one

@@
Typical infinite mixing automorphisms are rank-one
@@
Bashtanov proved that generic mixing automorphisms of probability space with respect to the Alpern-Tikhonov metric had rank one. Using Sidon’s constructions, we show that generic infinite mixing automorphisms also are rank-one.
@@ 1 Introduction
Alpern and Tikhonov [1],[7] introduced a complete metric on the space of mixing actions. Bashtanov [2] proved that the class of rank-one actions is generic in the space ${\bf Mix}$ of invertible mixing transformations with invariant probability measure. We will establish the same for the space ${\bf Mix_{\infty}}=\{T:\,T^{n}\to_{w}0\}$ of transformations of $T$ with sigma-finite invariant measure. We prove that the conjugacy class of the Sidon transformation of rank one (see [5], [6]) is dense in ${\bf Mix_{\infty}}$. Since the class of rank-one transformations is a $G_{\delta}$-set, taking into account the density of the latter, we obtain the genericity of this class. A consequence of this and the result of [3] is the typicalness of the minimality of the centralizer (the tipical transformation $T$ commutes only with its powers).
@@
where $\{A_{i}\}$ is some fixed family dense in the family of all measurable sets of finite measure. The space $({\bf Aut_{\infty}},\rho)$ is a complete separable metric space. It is easy to show that in $({\bf Aut_{\infty}},\rho)$ the so-called mixing transformations form a set of the first category (countable intersection of nowhere dense sets). If we strengthen the Halmos metric to the Alpern-Tikhonov metric, we obtain a complete metric space ${\bf Mix}_{\infty}$. Which invariants (i.e. properties preserved under conjugation in $\bf Aut_{\infty}$) are typical? A property is called typical if a dense $G_{\delta}$-set of transformations has this property.
@@
Fundamental in $({\bf Mix_{\infty}},r)$ a sequence $T_{k}$ converges to some $T$ by virtue of the completeness of the Halmos metric in the space ${\bf Aut}(\mu_{\infty})$. It is easy to show that this $T_{k}\to T$ converges to $T$ in the metric $r$ as well.
From the known methods, it can be shown that
for generic transformations in ${\bf Mix}_{\infty}$ that spectra of their symmetric powers are simple.
It follows from the results of [6] that
a typical transformation in ${\bf Mix}_{\infty}$ is not isomorphic to its inverse.
2 Density of the conjugacy classes of the Sidon automorphisms
Let us recall necessary definitions.
Construction of rank one. Fix a natural number $h_{1}$, a sequence $r_{j}\to\infty$ (parameter $r_{j}$ is the number of columns into which the tower of stage $j$ is cut) and a sequence of integer vectors (parameters of spacers)
$\bar{s}_{j}=(s_{j}(1),s_{j}(2),\dots,s_{j}(r_{j}-1),s_{j}(r_{j})),\ s_{(}i)\geq 0.$
At step $j=1$ a set of disjoint half-intervals $E_{1},SE_{1},S^{2}E_{1},\dots,S^{h_{1}-1}E_{1}$ of the same measure is given.
Let at step $j$ a system of non-intersecting half-intervals
@@ 2 Density of the conjugacy classes of the Sidon automorphisms
be defined, and on the half-intervals $E_{j},SE_{j},\dots,S^{h_{j}-2}E_{j}$ the transformation $S$ is the parallel translation. Such a set of half-intervals is called the tower of step $j$, their union is denoted by $X_{j}$ and is also called a tower.
@@
Sidon constructions. Let a rank one construction $S$ have the following property: the intersection $X_{j}\cap S^{m}X_{j}$ for $h_{j}<m\leq h_{j+1}$ can be contained only in one of the columns $X_{i,j}$ of the tower $X_{j}$. Such a construction is called Sidon. The measure of its space $X$ is infinite. If $r_{j}\to\infty$, then Sidon $S$ is mixing.
Theorem 1. Let $T$ be a mixing transformation, $S$ a Sidon mixing transformation of rank one. For every $\varepsilon>0$ there exists a conjugation $R$ such that $r(R^{-1}SR,T)<\varepsilon$.
@@
The proof of the lemma is simple, let it be an exercise.
The proof of the theorem (in brief).
Let us consider a part of the Kakutani tower:
@@
We take a very thin tower $X_{j}$ (so very large $j$), appearing in the construction of the Sidon transformation $S$, and wind on $U_{N}$ the tower $\tilde{X}_{j}$ corresponding to the conjugation $\tilde{S}=R^{-1}SR$ in a way that the following conditions are satisfied:
@@
(2) after the moving floor of the wound tower leaves the top of $U_{N}$ for the $k$-th time, it must travel outside $U_{N}$ for a time $s_{k}>10s_{k-1}$ ($s_{1}>10N$). This Sidon winding will ensure that the intersection measure is small:
@@
In reality, for a large $N$ and a suitable Sidon winding, instead of equalities we get inequalities with a very small right-hand side.
So for $\varepsilon>0$ there are $j$, $N$ and $R$ such that at the intersection $\tilde{X}_{j}\cap U_{N}$ the transformations $\tilde{S}$ and $T$ almost coincide, hence, $\tilde{S}$ and $T$ are close in the Halmos metric, but their powers are uniformly close with respect to the $d_{w}$ metric. Thus we obtain $r(R^{-1}SR,T)<\varepsilon$ for a suitable $R$.
The set of rank-one transformations is a $G_{\delta}$ set (use the proof from [2]). We have found that the class of mixing rank one transfomations is dense in $\bf Mix_{\infty}$. Thus, it is typical.
In [4] we used a similar trick, but the windings were random Ornstein and the spacers for flows and automorphisms of rank one were random Ornstein as well.
In the case of infinite measure, a pleasant opportunity arises to use large Sidon spacers, which gives us a simpler and constructive method. For other results on Sidon transformations, see in [5].
@@
Альперн и Тихонов [1],[7] ввели полную метрику на пространстве перемешивающих действий. Баштанов [2] доказал, что класс действий ранга один типичен в пространстве ${\bf Mix}$братимых перемешивающих преобразований с инвариантной вероятностной мерой. Мы установим аналогичную теорему для пространства ${\bf Mix_{\infty}}=\{T:\,T^{n}\to 0\}$реобразований $T$игма-конечной инвариантной мерой. Докажем, что класс сопряженности сидоновского преобразования ранга один (см. [5], [6]) плотен в ${\bf Mix_{\infty}}$. Так как класс преобразований ранга один является $G_{\delta}$-множеством, с учетом плотности последнего получаем типичность этого класса. Следствием этой типичности с учетом результата [3] является типичность минимальности централизатора (преобразование $T$оммутирует только со своими степенями).
@@
  • [1] Alpern St., Conjecture: In general a mixing transformation is not two-fold mixing, Ann. Probab., 13:1 (1985), 310-313